# Grid-based Multi-Robot Coverage Path Planning

### v1

@@ V The LS-MCPP Framework
Figure 7: The flowchart illustrates our LS-MCPP algorithmic framework on instance floor-small. Given an initial MCPP solution $\Pi_{0}$, LS-MCPP iteratively samples a pool $O$ from three pools of dedicated operators and then samples an operator $o$ from $O$, to transform a set $\mathcal{G}$ of subgraphs from one to another. It then applies ESTC on the subgraphs to evaluate the new MCPP solution $\Pi$, which is adaptively accepted according to its makespan $\tau$. A periodic subroutine of forced deduplication is called at the end of each iteration, and the best solution $\Pi^{*}$ is recorded if $\Pi$ is better with a lower makespan.
In this section, we solve MCPP defined in Problem LABEL:problem:_mcpp leveraging ESTC as a subroutine in our proposed local search algorithmic framework, namely LS-MCPP. Local search is a heuristic optimization method that directly explores the solution neighborhood to find a good solution to the problem. As depicted in Fig. 7, given an initial solution for MCPP instance $(G,I,R)$, its corresponding set $\mathcal{G}_{0}=\{G_{i}=(V_{i},E_{i})\}_{i=1}^{k}$ of $k$ connected subgraphs are computed. LS-MCPP operates as an iterative procedure utilizing a hierarchical sampling scheme to sample an operator to progressively transform the subgraph set $\mathcal{G}$ from one to another. It then employs ESTC on each subgraph $G_{i}$ to generate its coverage path and evaluate the makespan of MCPP. In the following subsections, we first provide an overview of LS-MCPP and then detailed descriptions of its key components.
V-A Algorithm Description
Overview: Consider the MCPP instasnce $(G,I,R)$, LS-MCPP incorporates three types of operators to alter a set $\mathcal{G}=\{G_{i}\}_{i=1}^{k}$ of subgraphs. We classify each subgraph $G_{i}$ as a light subgraph if its coverage path cost $c(\pi_{i})\leq\bar{c}$ by applying ESTC on each subgraph, where $\bar{c}$ is the average cost of all the coverage paths. Conversely, if its cost is greater than $\bar{c}$, we call $G_{i}$ a heavy subgraph. LS-MCPP utilizes the grow operators on light subgraphs to assign new vertices to cover and deduplicate operators on heavy subgraphs to eliminate unnecessarily repeated coverage. In addition, it employs exchange operators to balance the costs between subgraphs with a high cost difference. The operators are introduced in Sec. V-B, whose aims are to attain a cost-equilibrium MCPP solution with a low makespan. In general, LS-MCPP adopts a two-level hierarchical sampling approach to efficiently explore the constructed neighborhood in an iterative fashion. The first layer selects an operator pool using roulette wheel selection [goldberg1989genetic] from three pools, each containing a specific type of operator. The second layer heuristically draws an operator from the pool selected in the first layer, as detailed in Sec. V-C.
@@ 8       $O\sim\mathcal{O}$triangleright sample a pool by the probability $\sigma(\mathbf{p})$
9       update the weight $\mathbf{p}[O]$ of the selected pool $O$ to $(1-\gamma)\cdot\mathbf{p}[O]+\gamma\cdot\max(-\Delta\tau,0)$
@@ 11       $o\sim O$triangleright sample an operator by probability $\sigma(\mathbf{h})$
12       $\mathcal{G}^{\prime}\leftarrow$ apply operator $o$ on set $\mathcal{G}$
13       $\Pi^{\prime}\leftarrow$ apply ESTC on each subgraph in $\mathcal{G}^{\prime}$
14       if $\tau(\Pi^{\prime})<\tau(\Pi)$ then
@@ 13       $\Pi^{\prime}\leftarrow$ apply ESTC on each subgraph in $\mathcal{G}^{\prime}$
18             update $\mathcal{G}$ to $\mathcal{G}^{\prime}$ and $\Pi$ to $\Pi^{\prime}$ with a probability of $\exp{(\frac{\tau(\Pi)-\tau(\Pi^{\prime})}{t})}$
@@
21       if $i\,\%\,S=0\;\,\text{or}\;\,\tau(\Pi^{\prime})<\tau(\Pi)$ then
@@
Pseudocode (Alg. 2): LS-MCPP begins by computing the corresponding set $\mathcal{G}_{0}$ of subgraphs on the input initial MCPP solution $\Pi_{0}$ [Line 2]. It then initializes a set of pools via a subroutine Update-Pools described in Sec. V-B, each containing dedicated operators of the same type [Lines 2-2]. The pool weight vector $\mathbf{p}$ for the pools and the temperature scalar $t$ are also initialized [Lines 2-2]. The scalar $t$ dynamically adjusts in each iteration [Line 2] to make LS-MCPP progressively harder to accept a non-improving operator over iterations, as per the simulated annealing strategy [van1987simulated] to escape from local minimums. During each iteration [Line 2], vector $\mathbf{p}$ is utilized for roulette wheel selection [goldberg1989genetic] to select the most promising pool $O$ from $\mathcal{O}$ [Line 2], where $\sigma(\mathbf{p})$ denotes the softmax function of $\mathbf{p}$. After selecting pool $O$, its corresponding weight $\mathbf{p}[O]$ is updated with the specified weight decay factor $\gamma$ [Line 2]. Similar to the pool selection, LS-MCPP samples an operator $o$ from the selected pool $O$ with $\sigma(\mathbf{h})$ as the probability function [Line 2]. Here, $\mathbf{h}$ represents the vector of heuristic function $h$ evaluated on all operators in $O$ [Line 2]. Following this, LS-MCPP obtains a new set $\mathcal{G}^{\prime}$ by applying $o$ on the current set $\mathcal{G}$ of subgraphs [Line 2], and then generates the corresponding solution $\Pi^{\prime}$ by applying ESTC on each subgraph in $\mathcal{G}^{\prime}$ [Line 2]. LS-MCPP immediately accepts $o$ with the new $\mathcal{G}^{\prime}$ and the new $\Pi^{\prime}$ if $\Pi^{\prime}$ is better, and otherwise accepts it with a probability of $\exp{(-\Delta\tau/t)}\in[0,1]$ [Lines 2-2]. Upon the acceptance of $o$, LS-MCPP updates the set $\mathcal{O}$ of operator pools via Update-Pools [Line 2]. LS-MCPP also calls a subroutine Forced-Deduplication (described later) every $S$ iteration or when the current solution is better [Line 2]. The iterative search procedure terminates when it reaches the maximum number of iterations $M$, and returns the best solution $\Pi^{*}$ recorded [Line 2].
Forced Deduplication: A critical aspect contributing to the success of LS-MCPP is restricting duplication to only those necessary over the neighborhood exploration, such as those in a narrow path connecting two separate regions of graph $G$. For this purpose, LS-MCPP uses the Forced-Deduplication subroutine as described in Alg. 3 to deduplicate all possible duplication in two folds. The first part iterates through each path $\pi_{i}\in\Pi$ in a cost-decreasing order to remove any U-turns [Lines 3-3]. A U-turn of $\pi_{i}$ is defined as an edge $(u,v)\in E(\pi_{i})$ where $\exists\,(p,q)\in\pi_{i}$ such that $(u,p),(v,q)\in\pi_{i}$, thus forming a “U-turn” pattern of $p\rightarrow u\rightarrow v\rightarrow q$. It follows that we can safely remove $u,v$ to reduce $c(\pi)$ without introducing any new duplication or losing any vertex coverage if $u,v\in V_{i}\cap V^{+}$. Fig. 8 illustrates the procedure to remove U-turns for a path segment progressively. A special case of U-turn often occurs in incomplete hypervertices when $p\rightarrow u\rightarrow v\rightarrow q$ collapses into one edge if $p=v$, where we can safely remove $u,v$ without requiring $v\in V_{i}\cap V^{+}$. The second part [Lines 3-3] recursively applies every deduplicate operator from $O_{d}$ with two prioritizations of (1) higher cost of the corresponding path $\pi_{i}$ and (2) larger heuristic value of the operator. After the above two procedures, it updates all the pools in $\mathcal{O}$ by calling Update-Pools subroutine [Line 3]. It is worth noting that the first part can remove edges that do not constitute deduplicate operators, and some deduplicate operators can remove edges that are not U-turns depending on the MST for ESTC. This complementary relation contributes to the improved performance of LS-MCPP.
@@
2for $i\in\operatorname*{arg\,sort}\left(\{c(\pi_{i})\}_{i=1}^{k}\right)$ do
@@
8for $i\in\operatorname*{arg\,sort}\left(\{c(\pi_{i})\}_{i=1}^{k}\right)$ do
9       $O\leftarrow\{o_{d}(j,e)\in O_{d}\,|\,i=j\}$
10       for $o\leftarrow$ operator from $O$ with largest heuristic do
@@
13
14call Update-Pools ($\mathcal{G}$, $\mathcal{O}$, $\Pi$)
@@
Initial Solution: For the choice of the initial solution $\Pi_{0}$ for LS-MCPP, we limit ourselves to existing multi-tree MCPP methods. In addition, the method used to generate $\Pi_{0}$ should also be time-efficient to produce a MCPP solution with good quality. To this end, we choose our initial solution between the solutions resulting from VOR and MFC depending on whose quality is better. Both are integrated with the proposed ESTC algorithm described in Sec. IV-D.
@@
Our LS-MCPP framework is inherently better suited to utilizing a multi-tree method for generating an initial MCPP solution, which aligns with its design to iteratively improve a set of $k$ subgraphs using a local search strategy and apply ESTC to obtain coverage paths for evaluating the solution cost in each iteration. While a single-tree method can also provide initial subgraphs for LS-MCPP, induced by its solution coverage paths, our empirical finding [tang2024large] indicates that this approach typically results in significant overlaps among subgraphs and tends to reduce the effectiveness of the neighborhood operators within our LS-MCPP framework.
V-B Pools of Operators
We introduce three types of operators: grow operators, deduplicate operators, and exchange operators. These operators are devised to modify the boundaries of each connected subgraph $G_{i}=(V_{i},E_{i})\in\mathcal{G}$ of the input graph $G=(V,E)$. To start with, we define the duplication set $V^{+}=\{v\in V\,|\,n_{v}>1\}$ as the set of vertices covered by more than one path, where $n_{v}=\sum_{i\in I}\left|\{x\in V_{i}|x=v\}\right|$ counts the occurrences of the vertex $v\in V$ in all subgraphs. An operator alters the set $\mathcal{G}$ while ensuring that its property of complete coverage (i.e., $\bigcup_{i\in I}V_{i}=V$) and subgraph connectivity remain invariant. Each operator aims to introduce only necessary duplication, which has the potential to enhance the MCPP solution. Specifically, we consider edge-wise and vertex-wise operations for the above three types of operators. The vertex-wise operator only alters a single vertex one at a time, whereas the edge-wise operator alters a single edge one at a time. In practice, we find that altering an edge $(u,v)\in E$ that connect vertices belonging to different hypervertices (namely $\delta_{u}\neq\delta_{v}$) often brings unnecessary duplications on vertex coverage. Therefore, we restrict the edge-wise operators to only alter edges from a subset of $E$, given as $F=\{(u,v)\in E\,|\,\delta_{u}=\delta_{v}\}\subset E$.
Grow Operator: Let $B_{i}$ denote the set of vertices adjacent to but not part of $V_{i}$, defined as $B_{i}=\{v\in V_{i}\,|\,\exists(u,v)\in E,u\in V\setminus V_{i}\}$. A vertex-wise grow operator $o_{g}^{\text{(v)}}(i,\{v\})$ adds vertex $v\in B_{i}$ and all its adjacent edges to subgraph $G_{i}$. An edge-wise grow operator $o_{g}^{\text{(e)}}(i,\{u,v\})$ expands subgraph $G_{i}$ to add an edge $e=(u,v)\in E_{i}\cap F$ and all its adjacent edges into $G_{i}$, where vertices $u,v\in B_{i}$. We say $o_{g}(i,e)$ is valid iff $\exists(p,q)\in E_{i},(u,p),(v,q)\in E_{i}$. In essence, a valid $o_{g}(i,e)$ can only add an edge $(u,v)$ if there exists a parallel edge $(p,q)$ in $G_{i}$, as shown in the two examples in Fig. 9-(b)(c). This design choice circumvents the introduction of undesirable routing to the coverage path, as illustrated in Fig. 9-(a).
@@ V-B Pools of Operators
Deduplicate Operator: The deduplicate operator aims to remove unnecessary duplications from a subgraph. For vertex-wise operator $o_{d}^{\text{(v)}}(i,\{v\})$, it removes vertex $v\in V^{+}$ and all its adjacent edges from $G_{i}$. For edge-wise operator $o_{d}^{\text{(e)}}(i,\{u,v\})$, it removes edge $(u,v)\in E_{i}\cap F$ and all its adjacent edges from $G_{i}$, where $u,v\in V_{i}\cap V^{+}$ are already covered by other subgraphs. For edge-wise operator $o_{d}^{\text{(e)}}(i,\{u,v\})$, it is always valid if $\delta_{u}$ contains only $u$ and $v$, as removing them always reduces the cost of the corresponding ESTC path of $G_{i}$ (see Fig. 10-(c)). When $\delta_{u}$ contains not only $u$ and $v$, we say $o_{d}^{\text{(e)}}(i,\{u,v\})$ is valid iff it satisfies the following conditions: (1) any vertex contracted to $\delta_{e}^{t}$ is not in $V_{i}$; (2) all the four vertices contracted to $\delta^{b}_{e}$ are in $V_{i}$; (3) if there exist some $v\in V_{i}$ that is contracted to $\delta_{e}^{l}$ (or $\delta_{e}^{r}$), then all the four vertices contacted to $\delta_{e}^{l}$ (or $\delta_{e}^{r}$) are in $V_{i}$, and all the four vertices contacted to the (only) common neighbor between $\delta_{e}^{l}$ (or $\delta_{e}^{r}$) and $\delta_{e}^{b}$ are in $V_{i}$. The hypervertices $\delta^{t}_{e}$, $\delta^{b}_{e}$, $\delta^{l}_{e}$, and $\delta^{r}_{e}$ denote the four neighbors of $\delta_{u}$ (also equal to $\delta_{v}$) based on the position of $u$ and $v$ within $\delta_{u}$, as shown in Fig. 10-(a) that can be rotated from various symmetric cases. Such validity filtering on the edge-wise deduplicate operators is essential to avoid introducing unnecessary duplications to the ESTC path of the subgraph (see Fig. 10-(b)(c)(d)).
@@
Exchange Operator: Intuitively speaking, an exchange operator combines a grow operator and a deduplicate operator so that the deduplicate operator immediately removes the new duplication introduced by the grow operator, potentially improving the effectiveness of a single operator and speeding up the convergence of the local search. The vertex-wise operator $o_{e}^{\text{(v)}}(i,j,\{v\})$ is constructed from $o_{g}^{\text{(v)}}(i,\{v\})$ and $o_{d}^{\text{(v)}}(j,\{v\})$, where $v\in B_{i}\cap V_{j}$. The edge-wise operator $o_{e}^{\text{(e)}}(i,j,\{u,v\})$ is constructed from $o_{g}^{\text{(e)}}(i,\{u,v\})$ and $o_{d}^{\text{(e)}}(j,\{u,v\})$, Note that the deduplicate operator being combined might not be valid as $v\notin V^{+}$ for vertex-wise ones (or $u,v\notin V^{+}$ for edge-wise ones) are not necessarily covered by other subgraphs before the grow operator is applied. In practice, when updating $O_{e}$ (line 4 of Alg. 4), we iterate each grow operator $o_{g}(i,V_{o_{g}})$ and check if there exists deduplicate operator $o_{d}(j,V_{o_{d}})$ such that $V_{o_{d}}=V_{o_{g}}$ without requiring $V_{o_{d}}\subseteq V^{+}$.
Update Pools: In practice, we find that edge-wise operators work well most of the time but are sometimes unable to be constructed when encountered with many incomplete hypervertices, resulting in early termination of the LS-MCPP search. Vertex-wise operators are less restricted than edge-wise operators, however, tend to introduce unnecessary vertex duplications if applied near complete hypervertices. Therefore, we consider the edge-wise operators primary and make vertex-wise complementary when no edge-wise operators can be found. This is reflected in the Update-Pools function in Alg. 4 to initialize the set $\mathcal{O}$ of pools of operators [Line 4] or update the set $\mathcal{O}$ whenever $\mathcal{G}$ is changed by some operator $o$ [Lines 4-4]. Given the set $\Delta V$ of dirty vertices (defined in line 4 or line 4), Update-Pools first removes all operators involving any $v\in\Delta V$ from each pool [Line 4]. For each $v\in\Delta V$, it then iterates each light subgraph to add a valid edge-wise grow operator to $O_{g}$ if one exists; otherwise, it adds a vertex-wise grow operator [Lines 4-4]. Similarly, deduplicate operators are added to $O_{d}$ by iterating each heavy subgraph [Lines 4-4]. Finally, the pool $O_{e}$ of exchange operators is updated by combining the newly added operators in $O_{g}$ and $O_{d}$ [Line 4].

### v2

@@ V The LS-MCPP Framework
In this section, we introduce LS-MCPP, our local search algorithmic framework designed to effectively address MCPP as formulated in Problem 1. This framework integrates three types of neighborhood operators and utilizes ESTC as a subroutine, enhancing its ability to solve MCPP even in complex environments with incomplete hypervertices. LS-MCPP aims to optimize coverage paths by iteratively refining subgraphs and their corresponding coverage paths, ensuring both efficiency and completeness in coverage solutions.
V-A Overview
As shown in Fig. (7), LS-MCPP begins with an initial solution for the MCPP instance $(G,I,R)$ with its corresponding set $\mathcal{G}_{0}=\{G_{i}=(V_{i},E_{i})\}_{i=1}^{k}$ of $k$ connected subgraphs induced by the solution paths. LS-MCPP operates as an iterative procedure that employs a two-layer hierarchical sampling scheme, progressively adjusting the subgraph set $\mathcal{G}$ to efficiently explore the constructed neighborhood: The first layer selects an operator pool using roulette wheel selection [58] from three operator pools, each containing a different type of operators; The second layer heuristically selects an operator from the chosen pool, as detailed in Sec. V-D.
We categorize $G_{i}$ as a light subgraph if its coverage path cost $c(\pi_{i})$ from ESTC is no larger than the average coverage path cost $\bar{c}$ over all subgraphs; otherwise, it is a heavy subgraph. We define the duplication set $V^{+}=\{v\in V\,|\,n_{v}>1\}$ as the set of vertices included in more than one subgraph, where $n_{v}=\sum_{i=1}^{k}\left|\{x\in V_{i}|x=v\}\right|$ counts the occurrences of vertex $v\in V$ in all subgraphs.
LS-MCPP employs grow operators on light subgraphs to assign them new vertices to cover, deduplicate operators on heavy subgraphs to eliminate unnecessary duplication, and exchange operators to balance path costs between subgraphs with a large cost difference. The operators are detailed in Sec. V-C, designed to attain a cost-equilibrium MCPP solution with a low makespan. LS-MCPP then employs ESTC on each subgraph $G_{i}$ to generate its coverage path and evaluate the makespan. LS-MCPP also periodically calls a deduplication function to exploit the current subgraphs to achieve a low-makespan solution.
V-B Algorithm Description
@@ 8       $O\sim\mathcal{O}$triangleright sample a pool by the probability $\sigma(\mathbf{p})$
9       update the weight $\mathbf{p}[O]$ of the selected pool $O$ to $(1-\gamma)\cdot\mathbf{p}[O]+\gamma\cdot\max(-\Delta\theta,0)$
@@ 11       $o\sim O$triangleright sample an operator by probability $\sigma(\mathbf{h})$
12       $\mathcal{G}^{\prime}\leftarrow$ apply $o$ to its relating subgraph(s) of $\mathcal{G}$
13       $\Pi^{\prime}\leftarrow$ apply ESTC on updated subgraph(s) in $\mathcal{G}^{\prime}$
14       if $\Delta\theta\leftarrow\theta(\Pi^{\prime})-\theta(\Pi)<0$ then
@@ 13       $\Pi^{\prime}\leftarrow$ apply ESTC on updated subgraph(s) in $\mathcal{G}^{\prime}$
18             update $\mathcal{G}$ to $\mathcal{G}^{\prime}$ and $\Pi$ to $\Pi^{\prime}$ with a probability of $\exp{(\frac{-\Delta\theta}{t})}$
@@
21       if $i\,\%\,S=0\;\,\text{or}\;\,\Delta\theta<0$ then
@@
Pseudocode (Alg. 2): LS-MCPP begins by computing the corresponding set $\mathcal{G}_{0}$ of subgraphs on the input initial MCPP solution $\Pi_{0}$ [Line 2]. It then initializes a set of three operator pools via a subroutine Update-Pools described in Sec. V-C, containing only grow operators, deduplicate operators, and exchange operators, respectively [Lines 2-2]. It also initializes the temperature scalar $t$ and the pool weight vector $\mathbf{p}$ [Lines 2-2]. The scalar $t$ is scaled down by the decay factor $\alpha\in[0,1]$ for every iteration [Line 2] to dynamically adjust the probability for LS-MCPP to become increasingly less likely to accept a non-improving solution over iterations, which is the simulated annealing strategy [59] to skip local minima. The vector $\mathbf{p}$ represents the pool weight for each of the three aforementioned pools. During each iteration [Line 2], vector $\mathbf{p}$ is used for roulette wheel selection [58] to select a pool $O$ from $\mathcal{O}$ [Line 2], where $\sigma(\mathbf{p})$ is the softmax function of $\mathbf{p}$ determining the probability of selecting each pool. Once the pool $O$ is selected, its corresponding weight $\mathbf{p}[O]$ is updated with a weight decay factor $\gamma\in[0,1]$ [Line 2]. Similar to the pool selection, LS-MCPP samples an operator $o$ from the selected pool $O$ with $\sigma(\mathbf{h})$ as the probability function [Line 2]. Here, $\mathbf{h}$ is the vector of heuristic function $h$ evaluated on all operators in $O$ [Line 2], detailed in Sec. V-D. Following this, LS-MCPP applies operator $o$ to its relating subgraph(s) to obtain a new set $\mathcal{G}^{\prime}$ of subgraphs and a new solution $\Pi^{\prime}$ [Lines 2-2] and calculates the makespan increment $\Delta\theta$ to determine whether to accept the new solution $\Pi^{\prime}$. LS-MCPP accepts the new solution $\Pi^{\prime}$ if it has a smaller makespan, and, otherwise, accepts it with a probability of $\exp{(-\Delta\theta/t)}\in[0,1]$ [Lines 2-2]. Upon the acceptance of $\Pi^{\prime}$, LS-MCPP updates the set $\mathcal{O}$ of operator pools via Update-Pools (Alg. 4) [Line 2]. LS-MCPP also calls a subroutine Forced-Deduplication (Alg. 3) every $S$ iteration or the makespan has been reduced in the current iteration [Line 2]. The iterative search procedure terminates when it reaches the maximum number of iterations $M$, and LS-MCPP returns the best solution $\Pi^{*}$ found [Line 2].
@@
2for $i\in I$ sorted by descending $c(\pi_{i})$ do
@@
8for $i\in I$ sorted by descending $c(\pi_{i})$ do
9       for $o\in$ $O_{d}$ sorted by descending $h(o)$ do
@@
13call Update-Pools ($\mathcal{G}$, $\mathcal{O}$, $\Pi$)
@@
Forced Deduplication: A critical aspect contributing to the success of LS-MCPP is restricting duplication to only those necessary over the neighborhood exploration, such as those in a narrow path connecting two separate regions of graph $G$. For this purpose, LS-MCPP uses the Forced-Deduplication subroutine as described in Alg. 3 to deduplicate all possible duplication in two folds. The first part iterates through each path $\pi_{i}\in\Pi$ in a cost-decreasing order to remove any U-turns [Lines 3-3]. A U-turn of $\pi_{i}$ is defined as an edge traversal $(u\rightarrow v)$ in a “U-turn” pattern segment $p\rightarrow u\rightarrow v\rightarrow q$ in path $\pi_{i}$. If $u,v\in V_{i}\cap V^{+}$, then we can safely remove $(u\rightarrow v)$ to reduce $c(\pi)$ without losing any vertex coverage or introducing any new duplication. Fig. (8) illustrates the procedure to remove U-turns for a path segment progressively. A special case of U-turn often occurs for incomplete hypervertices where the “U-turn” pattern becomes $p=v\rightarrow u\rightarrow v\rightarrow q$. The second part [Lines 3-3] iteratively process each graph $G_{i}$ in descending order of path cost $\pi_{i}$ [Line 3], applying each deduplicate operator $o$ from $O_{d}$ in descending order of its heuristic value $h(o)$ [Line 3]. After the above two parts, the function re-initializes all the pools in $\mathcal{O}$ by calling Update-Pools (Alg. 4) [Line 3]. It is worth noting that the first part can remove edges that do not constitute deduplicate operators, and some deduplicate operators can remove edges that are not U-turns depending on the MST generated by ESTC. This complementary relation between the two parts contributes to the improved performance of LS-MCPP.
Initial Solution: For the choice of the initial solution $\Pi_{0}$ in LS-MCPP, we limit ourselves to existing multi-tree MCPP methods that can generate high-quality solutions efficiently. Specifically, we choose our initial solution from VOR or MFC, depending on which provides the better solution quality. Both methods are integrated with the proposed ESTC paradigm, as described in Sec. IV-D.
@@
Our LS-MCPP framework is inherently better suited to utilizing a multi-tree method for generating an initial MCPP solution, which aligns with its design to iteratively improve a set of $k$ subgraphs using a local search strategy and apply ESTC to obtain coverage paths for evaluating the solution cost in each iteration. While a single-tree method can also provide initial subgraphs for LS-MCPP, induced by its solution coverage paths, our empirical finding [21] indicates that this approach typically results in significant overlaps among subgraphs and tends to reduce the effectiveness of the neighborhood operators within our LS-MCPP framework.
V-C Pools of Operators
We introduce three types of operators: grow operators, deduplicate operators, and exchange operators. These operators are designed to modify the boundaries of each connected subgraph $G_{i}=(V_{i},E_{i})\in\mathcal{G}$ of the input graph $G=(V,E)$. An operator alters the set $\mathcal{G}$ while ensuring that its property of complete coverage (i.e., $\bigcup_{i=1}^{k}V_{i}=V$) and subgraph connectivity remains invariant. Each operator aims to introduce only essential duplication, with the potential to enhance solution quality.
@@ V-C Pools of Operators
Specifically, we consider both edge-wise and vertex-wise operations for these operators. An edge-wise operator alters a single edge at a time, whereas a vertex-wise operator alters a single vertex at a time. Altering a crossing edge $(u,v)\in E$ with $\delta_{u}\neq\delta_{v}$ often introduces unnecessary vertex duplications. Therefore, we restrict the edge-wise operators to only altering intra-hypervertex edges $E^{\text{intra}}=\{(u,v)\in E\,|\,\delta_{u}=\delta_{v}\}$. In practice, edge-wise operators generally perform well but sometimes cannot be constructed in configurations with many incomplete hypervertices, which may lead to early termination of the local search (Alg. 2). Vertex-wise operators are less restricted than edge-wise operators and tend to introduce unnecessary vertex duplications if applied to complete hypervertices. Therefore, vertex-wise operators serve only as complementary tools when no edge-wise operators are available, as reflected in the Update-Pools function (Alg. 4), which initializes or updates the operator pools. [^3] Our preliminary conference version [21] uses only edge-wise operators, which may lead to early termination issues. See Sec. VII-B for a detailed ablation study on the impact of vertex-wise operators.
Grow Operator: A grow operator serves to expand a subgraph to cover additional vertices already covered by other subgraphs. Let $B_{i}$ denote the set of “boundary” vertices that are not part of $G_{i}$ but adjacent to a vertex of $G_{i}$, defined as $B_{i}=\{v\in V\setminus V_{i}\,|\,\exists(u,v)\in E,u\in V_{i}\}$. A vertex-wise grow operator $o_{g}^{\text{(v)}}(i,\{v\})$ adds vertex $v\in B_{i}$ and all its adjacent edges to subgraph $G_{i}$. An edge-wise grow operator $o_{g}^{\text{(e)}}(i,\{u,v\})$ adds edge $e=(u,v)\in E_{i}\cap E^{\text{intra}}$ with $u,v\in B_{i}$ and all its adjacent edges into $G_{i}$, where $\exists(p,q)\in E_{i}$ such that $(u,p),(v,q)\in E_{i}$. In essence, a valid $o_{g}(i,e)$ can only add an edge $(u,v)$ if there exists a parallel edge $(p,q)$ in $G_{i}$, as shown in the two examples in Fig. (9b) and (9c). This design choice avoids introducing undesirable routing to the coverage path, as shown in the example in Fig. (9a). After its execution, a grow operator $o_{g}^{\text{(v)}}(i,\{v\})$ (or $o_{g}^{\text{(e)}}(i,\{u,v\})$) effectively adds vertex $v$ (or vertices $u$ and $v$) to $V^{+}$.
@@
Deduplicate Operator: A deduplicate operator serves to remove unnecessary duplication from a subgraph. A vertex-wise deduplicate operator $o_{d}^{\text{(v)}}(i,\{v\})$ removes vertex $v\in V_{i}\cap V^{+}$ and all its adjacent edges from $G_{i}$ such that $G_{i}$ remains connected. For an edge $e=(u,v)\in E_{i}\cap E^{\text{intra}}$, let $\delta^{\,t}_{e}$, $\delta^{\,b}_{e}$, $\delta^{\,l}_{e}$, and $\delta^{\,r}_{e}$ denote the four neighbors of $\delta_{u}$ (also equal to $\delta_{v}$) based on the positioning of $u$ and $v$. Fig. (10a) demonstrates one example that can be rotated to various symmetric cases (with only $\delta^{\,t}_{e}$ potentially containing both a vertex adjacent to $u$ and one adjacent to $v$). An edge-wise deduplicate operator $o_{d}^{\text{(e)}}(i,\{u,v\})$ removes edge $(u,v)\in E_{i}\cap E^{\text{intra}}$ with $u,v\in V_{i}\cap V^{+}$ and all its adjacent edges from $G_{i}$ such that $G_{i}$ remains connected. If $\delta_{u}$ contains only $u$ and $v$, then this removal is always valid and reduces the cost of the corresponding ESTC path of $G_{i}$, as shown in Fig. (10c). Otherwise, this removal is constrained by three conditions: (1) any vertex contracted to $\delta_{e}^{t}$ is not in $V_{i}$; (2) all the four vertices contracted to $\delta^{b}_{e}$ are in $V_{i}$; and (3) if any $v\in V_{i}$ is contracted to $\delta_{e}^{l}$ (or $\delta_{e}^{r}$), then both $\delta_{e}^{l}$ (or $\delta_{e}^{r}$) and its (only) common neighboring hypervertex with $\delta_{e}^{b}$ must be complete, and all their vertices must also be in $V_{i}$. Such validity filtering on the edge-wise deduplicate operators is essential to avoid introducing unnecessary duplications to the corresponding ESTC path of $G_{i}$, as shown in Fig. (10). After its execution, a duplicate operator $o_{d}^{\text{(v)}}(i,\{v\})$ (or $o_{d}^{\text{(e)}}(i,\{u,v\})$) effectively removes vertex $v$ (or vertices $u$ and $v$) from $V^{+}$ if the vertex (or vertices) is not included in multiple subgraphs.
Exchange Operator: An exchange operator combines a grow operator and a deduplicate operator so that the deduplicate operator immediately removes the new duplication introduced by the grow operator, potentially improving the effectiveness of a single operator and speeding up the convergence of the local search. A vertex-wise exchange operator $o_{e}^{\text{(v)}}(i,j,\{v\})$ combines $o_{g}^{\text{(v)}}(i,\{v\})$ and $o_{d}^{\text{(v)}}(j,\{v\})$, where $v\in B_{i}\cap V_{j}$. An edge-wise exchange operator $o_{e}^{\text{(e)}}(i,j,\{u,v\})$ combines $o_{g}^{\text{(e)}}(i,\{u,v\})$ and $o_{d}^{\text{(e)}}(j,\{u,v\})$. Note that the operators $o_{g}$ and $o_{d}$ being combined must be both valid except that vertex $v$ for a vertex-wise $o_{d}^{\text{(v)}}(j,\{v\})$ and vertices $u$ and $v$ for an edge-wise $o_{d}^{\text{(e)}}(j,\{u,v\})$ are not required to be in $V^{+}$ since they are not necessarily covered by subgraphs other than $G_{j}$ before the grow operator is applied. After the execution of an exchange operator, $V^{+}$ remains unchanged.
@@
Update Pools: The Update-Pools function in Alg. 4 initializes the set $\mathcal{O}$ of pools of operators [Line 4] or updates the set $\mathcal{O}$ whenever $\mathcal{G}$ is changed by some operator $o$ [Lines 4-4]. Given the set $\Delta V$ of dirty vertices (defined in line 4 or line 4), Update-Pools first removes all operators involving any $v\in\Delta V$ from each pool [Line 4]. For each $v\in\Delta V$, it then iterates each light subgraph to add a valid edge-wise grow operator to $O_{g}$ if one exists; otherwise, it adds a vertex-wise grow operator [Lines 4-4]. Similarly, deduplicate operators are added to $O_{d}$ by iterating each heavy subgraph [Lines 4-4]. Finally, the pool $O_{e}$ of exchange operators is updated by combining the newly added operators in $O_{g}$ and $O_{d}$ [Line 4].